# A Deep Implicit-Explicit Minimizing Movement Method for Option Pricing in Jump-Diffusion Models

### v1

@@ 5. Conclusions
This work develops a novel deep learning method for the pricing of European basket options in models that follow jump-diffusion processes. In order to address the intricacies of the problem, we introduce a decomposition technique, expressing the option price as the sum of an unknown component (time value) and a known lower-bound function (intrinsic value). The incorporation of a domain truncation method further enhances the accuracy of our numerical schemes. By projecting option prices onto a bounded subset and efficiently approximating solutions for extreme moneyness values within this truncated domain, we improve the robustness and reliability of our approximation method. Moreover, our research introduces two distinct approaches for computing the integral operator. We believe that the combination of the deep implicit-explicit minimizing movement methods, the decomposition of the solution, and the domain truncation method form a robust toolkit for enhancing the accuracy and efficiency of option pricing methods in advanced financial models.

### v2

@@ 5. Conclusions
This work develops a novel deep learning method for the pricing of European basket options in models that follow jump-diffusion processes. To address the intricacies of the problem, we introduce a decomposition technique, expressing the option price as the sum of an unknown component (time value) and a known lower-bound function (intrinsic value). The incorporation of a domain truncation method further enhances the accuracy of our numerical schemes. By projecting option prices onto a bounded subset and efficiently approximating solutions for extreme moneyness values within this truncated domain, we arrive at an accurate and reliable approximation of the underlying solution to the PIDE problem. The combination of the deep implicit-explicit minimizing movement methodology, the decomposition of the solution, and the domain truncation method form a robust toolkit for enhancing the accuracy and efficiency of option pricing methods in advanced financial models, by providing a complete solution at every point of the space-time domain. As such, it is straightforward to infer further quantities, such as Greeks, from the computed space-time solution. A basic comparison with popular and successful approaches further showcases the practical relevance of the proposed method.